Solving for Rates What annual rate of return is earned on a $2,000 investment when it grows to $4,500 in eleven years?

Present Value of an Annuity Due If the present value of an ordinary, 5-year annuity is $6,600 and interest rates are 11 percent, what's the present value of the same annuity due?

Dividend Initiation and Stock Value A firm does not pay a dividend. It is expected to pay its first dividend of $0.80 per share in 2 years. This dividend will grow at 12 percent indefinitely. Using a 14 percent discount rate, compute the current value of this stock.

Average Return The past five monthly returns for PG Company are 1.80 percent, -2.6 percent, 4.80 percent, 4.3 percent, and 2.5 percent. What is the average monthly return?
Answer #1

a.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

4500=2000*(1+r/100)^11

(4500/2000)^(1/11)=(1+r/100)

(1+r/100)=1.0765

r=1.0765-1

=7.65%(Approx).

2.Present value of annuity due=Present value of annuity*(1+rate)

=6600*1.11

=$7326

3.Value after year 2=(D2*Growth rate)/(Discount rate-Growth rate)

=(0.8*1.12)/(0.14-0.12)

=44.8

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=0.8/1.14^2+44.8/1.14^2

=$35.09(Approx).

4.Average monthly return=Total return/Total time period

=(1.8-2.6+4.8+4.3+2.5)/5

=2.16%